Week 49

 

Machine coordinators Mike Lamont and Ralph Assmann

 

 

Day

Date

Beam

No beam

%

Monday

30-Nov

22

2

 

Tuesday

1-Dec

19

5

 

Wednesday

2-Dec

3

21

 

Thursday

3-Dec

24

0

 

Friday

4-Dec

17

7

 

Saturday

5-Dec

24

0

 

Sunday

6-Dec

21

3

 

Week49

Beam/Nobeam

130

38

77

Day

Short summary

Link to report from 8:30 meeting

Tuesday 8th December

 

8:30 Summary for December 7

Monday 7th December

Since 00:30: Beam off.
a) Loss of cryo maintain in sector 81 led to dump of stable beams.
b) Electrical perturbation at 04:25 during recovery. Loss of warm compressor in IR4. Magnet trips. Other effects on machine elements.
c) Recovery estimated between 22:00 and 24:00.

8:30 Summary pptx for December 6

Sunday 6th December

00:00-02:30: Four bunch setup.
a) 1.5e10 p in beam1 in 4 bunches. Fully protected machine. 25h beam lifetime. Correctly dumped.
b) At 01:32: Two beams with each 4 bunches stored. B1: 1.4e10. B2: 1.5e10.
c) Dump tests show normal loss patterns both for bunched and unbunched beams.
02:30-06:30: Injecting for stable beams. Finally successful around 6:30 after resolving a lifetime problem with beam 2. 4 on 4 circulating and colliding with reasonable lifetime for both beams.
06:55-11:00: STABLE BEAMS!
First time for LHC! 450 GeV collisions for the experiments.
11:00-22:30: End of fill, recovery.
a) Stable beam was dumped due to BCT server reboot.
b) Switch off of all LHC sectors at 14h15.
c) Access from 16:00 to 18:00.
d) Pre-cycle and recovery.
e) Injection.
22:30-00:30: STABLE BEAMS.
Back to stable beams and 450 GeV collisions with increased bunch intensities of about 6e9 protons per bunch.

 

Saturday 5th December

01:00-02:00: Work with Beam2.
a) LHCb spectrometer non-closure (0.8mm) corrected.
b) Then RF problems and no beam.
01:00-06:00: Work with Beam1.
a) Beta beat studies and correction: Dipole b2, IR8 and IR2. Improvement for most arcs in horizontal beta-beating for beam1. Clear big error in IP7 remains.
b) 100 beam dumps to verify change in dump kicker software.
07:00-09:30: Prepare and perform pre-cycle.
09:30-10:00: Prepare injection.
10:00-12:00: Beam2 only. Correct tunes and chromaticity. Bring on ALICE spectrometer. Establish reference. Interrupted by several interlocks from RF. ALICE spectrometer non-closure (0.35mm) corrected.
12:00-14:00: Golden orbit and IR steering.
a) Beam-beam separation minimized in all IR's.
b) ALICE solenoid switched on.
c) Golden orbit loaded.
d) Orbit in cleaning insertions corrected to Sunday reference (collimatio beam-based alignment).
e) Software interlock on dipole correctors activated (for STABLE beam mode).
14:00-16:00: Reestablishing collimation.
a) Collimator settings reloaded. No reoptimization required (very good machine and collimator reproducibility). Loss patterns OK. Injection OK (~70% injection efficiency).
b) Identified lifetime reduction from transverse damper. Lifetime with collimators in nominal position and transverse damper off reached >25h.
c) BLM beam loss spectra recorded with collimator scans.
d) Most collimator position interlocks activated (except injection and dump protection).
16:00-17:00: RF work.
16:00-20:00: Dump and injection protection.
a) TDI's set up.
b) TCDQ and IR6 TCSG set up.
c) Tests for asynchronous dumps. OK.
d) Position interlocks on 2 TDI, 2 TCDQ, 2 IR6 TCSG, 2 TCLIA, 2 TCLIB.
e) ALICE solenoid inverted.
20:00-23:00: Trip of B24.R1. Recovery, pre-cycle and injection set-up.
23:00-00:00: Reestablish reference. Final dump protection checks.

9:00 Summary pptx for December 4

Friday 4th December

07:30-12:00: Aperture studies continued. Stopped by beam-induced quench in cell 15R2 (injecting into orbit bump). Beam intensity was around 4e9 p.
12:00-13:00: Prepare access (advanced this compared to schedule).
13:00-17:00: Access and other work:
a) Access for undulator, TOTEM, wire scanners, LBDS, RF damper and QPS.
b) Intervention on PS magnets.
c) RF changed configuration: work with half the number of cavities and about the same total voltage (7 MV instead of 8 MV) to reduce power dissipation on the klystron collectors.
d) Further analysis of tune ripple observation.
17:00-19:00: Pre-cycle. No problems, except trips RSF2.A78.B1 and RSD2.A78.B2.
19:00-20:00: Injection set-up. Unable to turn RF power on for module M1B1 for beam 1.
20:00-21:00: Beam 2:
a) Inject and correct beam 2.
b) 20:30 -> For the first time two bunches in a single LHC beam (total intensity 8e9 p in beam 2).
c) Systematic checks for instrumentation and beam dump. All OK.
21:00-22:00: Beam 2:
a) Undulator successfully at 200A, no light seen yet.
b) Switching on LHCb spectrometer to nominal 2106.371 micro-rad = 5850A.
c) RF triggered beam dump. RF problem related to 18 kV distribution problem.
22:00-01:00: Work on RF. Injection set-up.

8:30 summary pptx for December 3

Thursday 3rd December

23:00-02:00: IR8 aperture study.
Looks OK for LHCb. To be analyzed.
02:00-05:00: Pre-cycle.
05:00-08:00: Beta beat studies.
Beam 1 correction had a positive effect. Need to understand IR7 results more carefully for Beam 2.All trims and tunes reset to initial settings.
08:00-13:30: TDI and LBDS studies.
a) Ring injection protection set up for beam2 (+-6.8sigma for TDI, TCLIA, TCLIB plus TDI angle). Protection verified with injection failure tests with beam. Over-injection mode tested: requires some BLM interlocks disabled for this mode.
b) Dumpline trajectories verified for beam 1 and beam 2. Dump failure tests with beam.
c) IR6 aperture scans with bumps for both beams. No sign of abnormal aperture limits.
13:30-15:00: Beam instrumentation checks.
Q/Q' and coupling measurements.Tune stability issue H plane: tune walks around at 0.001 level.
15:00-23:00: IR aperture checks.
a) CMS aperture scan OK.
b) ALICE aperture scan OK.
c) In ATLAS the V plane is OK (respectively similar to the other experiments). The H plane is 'different' from all other IRs.
ATLAS dumped with the BCM on both beams, while the machine BLMs never gave any significant signal (in the V plane the BLMs gave the usual signals): both beam1 and beam2. Went to relatively large amplitudes, even though less than in other IRs. Be careful not to draw hasty conclusions.
00:00-07:30: Separation bumps and crossing angle.
a) All bumps switched on individually. No problems (polarity etc). Non-closure and dispersion verified.
b) Luminosity scan knobs tested successfully.
c) Finally at 06:25 all separation bumps and crossing angles (nominal injection values) switched on and both beams injected with NO major losses. IP8 losses not related to IR setup (stayed unchanged). No sign of problems with cross-talk from MCBXs (nested IP correctors).
d) Orbit saved and bumps removed.

8:30 summary pptx for December 2

Wednesday 2nd December

01:23 18 kV drop - Meyrin site. LHC off - cryo OK. Several accesses completed by noon. QPS tests. Issues with recovery as filters are put onto electricity lines, tripping various circuits. Magnets switched off for filters in further sectors. ALICE solenoid also tripped. Beam back at 22:30. IR8 aperture studies.

8:30 summary pptx for December 1

Tuesday 1st December

Beam dump/RF synchronization tests in the morning. Access at midday. Beam back into the LHC at 18:00.

8:30 summary pptx for November 30

Monday 30th November

Alice, Atlas and CMS solenoids successfully brought on - orbit and coupling corrected. Aperture and kick response measurement overnight. Wirescanners successfully calibrated.

8:30 summary pptx for November 29